6 years ago bob was 4 times as old as Mary. In 4 years he will be twice as old as Mary. How old are they?
Found 2 solutions by 90887896, richard1234:
Answer by 90887896(4)
Answer by richard1234(7193)

Let b and m be Bob's and Mary's ages. Then,

b-6 = 4(m-6) --> b-6 = 4m - 24

b+4 = 2(m+4) --> b+4 = 2m + 8

This is essentially a system of equations that can be solved for b and m. Solving using any method (I used substitution), we get m = 11, b = 26, so Mary is 11 and Bob is 26.